Question 4 Evaluate the line integral:

(i) of 𝑇(π‘₯) = 4π‘₯^3 along the line segment from (βˆ’2,1) to (1,2).

(ii) where the curve 𝐢 is parameterized through π‘₯(𝑑) = cos 𝑑, 𝑦(𝑑) = sin 𝑑 and 𝑧 = 𝑑^2 with 𝑑 ∈ [0,2πœ‹] of ∫(π’šπ’…π’™ + π’™π’…π’š + 𝒛𝒅𝒛) 𝐢

(iii) ∫ 𝐅(π‘₯, 𝑦, 𝑧) β‹… ⅆ𝐫 𝐢 , where 𝐹(π‘₯, 𝑦, 𝑧) = (5𝑧^2 , 2π‘₯, π‘₯ + 2𝑦) and the curve 𝐢 is given by π‘₯ = 𝑑, 𝑦 = 𝑑^2, and 𝑧 = 𝑑^2 with 𝑑 ∈ [0,1]


1
Expert's answer
2021-11-16T16:04:17-0500

(I)

Given"T(X)=4x^3"

Along line segment (-2,1) to (1,2(

Line integral is given by

"\int T(X) dx=\int_{X=-2}^{X=1}4x^3dx\\=\frac{4x^4}{4}\int_{X=-2}^{X=1}\\=1-(-2)^4\\=1-16\\Txdx=-15"

(ii)

Given X(t)=cos t,y(t)=sin t, z(t)=t2

dx=-sin ft, Dy=cos ft, dz=2tdt

Required lines integral is given by

"\int ydx+xdy+zdz=\int sin\space tx-sin\space tdt+cos \space tx+2t^3dt"

"\int_{t=0}^{2n}(cos^2t-sin^2t+2t^3)dt\\

\int_{t=0}^{2n}(cos\space 2t+2t^3)dt\\=\frac{1}{2}(2n)^4"

The required value of line integral is "2n^4"


(iii)

"Given\\x=t, y=t^2,z=t^2"

Such that

"dx=DT,Dy=2tdt, dz=2tdt"

Now "\int f(X,y,z)dt=\int_{t=0}^{t=1}(5t^2+4t^2+2t^2+4t^3)dt\\=\frac{5t^3}{5}+\frac{4t^3}{3}+\frac{2t^3}{3}+\frac{4t^4}{4}\\=1+2+1\\=4"

Required value of line integral is 4